# How Beginning Programmers and Code LLMs (Mis)read Each Other

### v1

@@ 5.1. Evaluation Plan
We collected three types of data which lend themselves to qualitative analysis: (1) information about student experience and demographics, (2) free-response questions about future use of Charlie, and (3) semi-structured interview responses. We employed both inductive and deductive open coding towards consensus. Our aim was to identify common themes present in this specific dataset, rather than to develop a theory. Two researchers with previous qualitative experience conducted the analysis. We present selected quotes from the surveys and interviews throughout. Quotations have been lightly edited to remove speech errors and any identifiable information. Each participant’s quote is accompanied by a pseudonym assigned to them during data collection.
@@
5.2. Measures of Success
There are several ways to measure success when evaluating the natural language to code task. The success rate is the fraction of all attempts on which the model generates a working program. Therefore, a participant who takes several attempts to solve a problem will have a lower success rate than another who succeeds in one try. We might also ask whether a participant is ever able to solve a problem; we refer to this as the eventual success rate. This metric considers only the participant’s final attempt at each assigned problem.
Although success rates measure the correctness of the code that students saw during the experiment, LLM generation is non-deterministic. [^2] Greedy generation is significantly worse for coding tasks than non-deterministic generation (Chen et al., 2021). Therefore, studying success rates can be misleading: a participant may have just been lucky with a bad prompt or unlucky with a good prompt. For this reason, we also employ an alternative metric called pass@1, which accounts for non-deterministic generation (Chen et al., 2021). Since the debut of Codex, pass@1 has become the standard metric used to evaluate LLMs on the natural language to code task, including GPT-4 (OpenAI, 2023), Code Llama (Baptiste Rozière et al., 2023), and other models (Li et al., 2023; Fried et al., 2022; Nijkamp et al., 2022).
Given a natural language prompt, pass@1 (Chen et al., 2021) is an estimate of the probability that the LLM will generate working code in one attempt. In the LLM development literature, the accepted best practice for computing pass@1 is to query the LLM 200 times for the same prompt and test every generated program (Chen et al., 2021; Fried et al., 2022; Xu et al., 2022; Rozière et al., 2023b). Sampling 200 generations for all 2,000+ prompts is very expensive with the Codex API. Instead, we use a recently released open Code LLM called StarCoder (Li et al., 2023) that is nearly as capable as the Codex model on Python benchmarks. Pass@1 with StarCoder will be slightly lower than Codex success rates because of model differences. However, pass@1 is a more stable measure of whether a prompt will succeed than success rate. We use pass@1 for the bulk of our analyses.
5.3. Positionality
We are affiliated with the three institutions where this research was conducted; we range from undergraduate students to tenured faculty. We developed the problem lists, problem difficulty ratings, and other elements of the study design within a shared educational context. Some authors are course instructors for CS1. As described in § 4.3, significant care was taken to address power dynamics between participants and researchers. Some authors also contribute to the development and evaluation of open-source Code LLMs. Overall, the potential incentives for the research team are complex, as we approach this work as both educators and researchers. We aspire to a neutral perspective on Code LLMs, while attempting to center the student experience.
@@ 6. RQ1: Do Students Succeed at Prompting Code LLMs with Natural Language?
Figure 5. Basic measures of student success at the natural language to code task. Success rate is the fraction of all attempts by a participant that succeed. Eventual success rate is the fraction of last attempts at a problem by a participant that succeed. Pass@1 resamples the LLM several times to estimate the probability of success. We present these measures by institution. Figure 4(a) presents the means. Figure 4(b) and figure 4(c) show the distribution of (eventual) success rates. Eventual success rates are higher than success rates, which is to be expected: Figure 4(d) shows that many students make several attempts at a problem before an eventual success or give up.
@@ 6.1. Basic Findings
6.2. Do Participants Find the Task Challenging?
Question Scale Mean
How mentally demanding was the task? Very low-¿Very high 4
How hurried was the pace of the task? Very low-¿Very high 3.3
How successful were you? Perfect-¿Failure 3.6
Table 1. Mean NASA-TLX ratings (Hart and Staveland, 1988)
@@ 6.2. Do Participants Find the Task Challenging?
Raised Monolingual Not in English 27 0.22
@@
Raised Multilingual Not in English 2 0.20
@@ 6.3. Who Succeeds at the Task?
Using data from the post-survey, we analyze the relationship between pass@1 rates and previous knowledge, prior programming experience, and demographics (see Table 2 for a summary of demographics). We find only two statistically reliable differences:
@@ 7. RQ2: Where do student difficulties come from?
Having shown that students find it hard to prompt a Code LLM in natural language (§ 6), we explore why. In this section, we present quantitative and qualitative results that address RQ2: when students struggle with the task, where do the struggles come from? What are the most challenging aspects of the natural language to code task?
@@ 7.1. What aspects of the task do students say are hard?
Charlie Doesn’t Understand Me 89
@@
Issues with Study Platform 8
@@
The most commonly raised issues related to Charlie’s understanding of prompts (n=89); we divided these into subcodes. One of the most common of these was the sentiment that Charlie failed to understand good descriptions (n=23). For instance, redCoyote commented, “It was definitely difficult to have a concept of what you wanted written in your head, and then feel like you’re articulating it well, but having it not work properly.” Similarly, aquaLadybug reports feeling helpless when a good prompt didn’t succeed: “if I was saying it […] how I thought was the best way to say it, but it still wasn’t working, I had no idea where to go from there.”
@@
Another major theme was issues with the generated code. Many related to perceived bugs in the generated code or difficulty debugging (26%). Students also mentioned finding the model’s randomness frustrating (8%). khakiBee was alarmed to find that resubmitting the same prompt could generate different programs, commenting “You feel like you’ve made progress, and then because it did a different thing the next time, it’s like, what do I change? I’m trying to change what I give to the cow. And then that should change what the cow is doing. But if I’m not changing anything, why is that changing?” Some students also experienced the opposite issue: despite changing their descriptions, the model generated the same incorrect function repeatedly. purpleCarp commented, “Sometimes I changed my […] description and it just repeated the code the same. And it’s just very frustrating”. This highlights the difficulty of working with stochastic models: students expect the model output to be faithful to their descriptions.
@@ 7.2. Which Problems Do Students Say Are Hard?
We find that pass@1 and eventual success rates both vary by category (Table 4). We fit a binomial mixed-effects model to pass@1, with fixed effects of category, institution, and their interaction, and random effects of problem and participant. A statistically reliable difference in pass@1 was observed only for Sorting problems, which were the most challenging ($p$=0.045). We fit a statistical model to explore whether category difficulty varied by institution and found that participants from Oberlin had lower pass@1 in the Nested category compared to other students, but that the effect was not reliable ($p$=0.063).
@@ 7.4. What Do Students Do When They Encounter Unfamiliar Python?
Figure 6. An example problem, exp, which was rated differently by Oberlin and Wellesley students, likely due to the list comprehension.
Declaration, Prompt & Completion
@@
Question
Is this code you would write yourself?
Student Responses
Wellesley: Yes, Oberlin: No

### v2

@@ 5.1. Evaluation Plan
We collected three types of data which lend themselves to qualitative analysis: (1) information about student experience and demographics, (2) free-response questions about future use of Charlie, and (3) semi-structured interview responses. We employed both inductive and deductive open coding towards consensus. Our aim was to identify common themes present in this specific dataset, rather than to develop a theory. Two researchers with previous qualitative experience conducted the analysis; Section A.2 contains details of the coding methodology. We present selected quotes from the surveys and interviews throughout. Quotations have been lightly edited from the automatically generated transcripts. This includes addressing grammar/punctuation, removing speech errors or filler words, and avoiding the disclosure of any identifiable information. Each participant’s quote is accompanied by a pseudonym assigned to them during data collection.
@@
Figure 5. Basic measures of student success at the natural-language-to-code task. Success rate is the fraction of all attempts by a participant that succeed. Eventual success rate is the fraction of last attempts at a problem by a participant that succeed. Pass@1 resamples the LLM several times to estimate the probability of success. We present these measures by institution. Figure 5(a) presents the means. Figure 5(b) and figure 5(c) show the distribution of (eventual) success rates. Eventual success rates are higher than success rates, which is to be expected: Figure 5(d) shows that many students make several attempts at a problem before an eventual success or give up.
5.2. Measures of Success
There are several ways to measure success when evaluating the natural-language-to-code task. The success rate is the fraction of all attempts on which the model generates a working program. Therefore, a participant who takes several attempts to solve a problem will have a lower success rate than another who succeeds in one try. We might also ask whether a participant is ever able to solve a problem; we refer to this as the eventual success rate. This metric considers only the participant’s final attempt at each assigned problem. The eventual success rate metric is likely specific to this paper, as closely related work (Prather et al., 2023; Denny et al., 2023; Kazemitabaar et al., 2023) studies different notions of success or does not permit controlled, repeated interactions.
Although success rates measure the correctness of the code that students saw during the experiment, LLM generation is non-deterministic. [^4] Greedy generation is significantly worse for coding tasks than non-deterministic generation (Chen et al., 2021). Therefore, studying success rates can be misleading: a participant may have just been lucky with a bad prompt or unlucky with a good prompt. For this reason, we also employ an alternative metric called pass@1, which accounts for non-deterministic generation (Chen et al., 2021). Since the debut of Codex, pass@1 has become the standard metric used to evaluate LLMs on the natural-language-to-code task, including GPT-4 (OpenAI et al., 2023), Code Llama (Rozière et al., 2024), and other models (Li et al., 2023; Fried et al., 2023; Nijkamp et al., 2022).
Given a natural language prompt, pass@1 (Chen et al., 2021) is an estimate of the probability that the LLM will generate working code in one attempt. In the LLM development literature, the accepted best practice for computing pass@1 is to query the LLM 200 times for the same prompt and test every generated program (Chen et al., 2021; Fried et al., 2023; Xu et al., 2022; Rozière et al., 2024). Sampling 200 generations for all 2,000+ prompts generated as part of this study would be very expensive with the Codex API. Instead, we use a recently released open Code LLM called StarCoder (Li et al., 2023) that is nearly as capable as the Codex model on Python benchmarks. Pass@1 with StarCoder will be slightly lower than Codex success rates because of model differences. However, pass@1 is a more stable measure of whether a prompt will succeed than success rate. We use pass@1 for the bulk of our analyses.
5.3. Positionality
All authors were affiliated with the institutions from which participants were recruited (Oberlin, Wellesley, or Northeastern) at the time of the study; we range from undergraduate students to tenured faculty. We developed the problem lists, problem difficulty ratings, and other elements of the study design within a shared educational context. The last three authors are course instructors for CS1. As described in § 4.3, significant care was taken to address power dynamics between participants and researchers. Some authors also contribute to the development and evaluation of open-source Code LLMs. Overall, the potential incentives for the research team are complex, as we approach this work as both educators and researchers. We aspire to a neutral perspective on Code LLMs, while attempting to center the student experience.
@@ 6.1. Basic Findings
Raised Monolingual Not in English 27 0.20
@@
Raised Multilingual Not in English 2 0.22
@@
6.2. Do Participants Find the Task Challenging?
Abbreviated Question Scale (1 to 7) Mean
How mentally demanding was the task? Very low-¿Very high 4
How hurried or rushed was the pace of the task? Very low-¿Very high 3.3
How successful were you? Perfect-¿Failure 3.6
Table 2. Mean NASA-TLX ratings (Hart and Staveland, 1988).
@@ 6.3. Who Succeeds at the Task?
Using data from the post-survey, we analyze the relationship between pass@1 rates and previous knowledge, prior programming experience, and demographics (see Table 1 for a summary of demographics). We find only two statistically reliable differences (see Appendix, Table 11 for the full statistical analyses):
@@ 7. RQ2: Where do student difficulties come from?
Having shown that students find it hard to prompt a Code LLM in natural language (§ 6), we explore why. In this section, we present quantitative and qualitative results that address RQ2: when students struggle with the task, where do the struggles come from? What are the most challenging aspects of the natural-language to-code task?
@@ 7.1. What aspects of the task do students say are hard?
Charlie Doesn’t Understand Me 91
@@
Issues with Study Platform 10
@@
The most commonly raised issues related to Charlie’s understanding of prompts (n=91); we divided these into subcodes. One of the most common of these was the sentiment that Charlie failed to understand good descriptions (n=23). For instance, redCoyote commented, “It was definitely difficult to have a concept of what you wanted written in your head, and then feel like you’re articulating it well, but having it not work properly.” Similarly, aquaLadybug reports feeling helpless when a good prompt didn’t succeed: “if I was saying it […] how I thought […] is the best way to say it, but it still wasn’t working, I had no idea where to go from there.”
@@
Another major theme was issues with the generated code. Many commments related to perceived bugs in the generated code or difficulty debugging (26%). Students also mentioned finding the model’s randomness frustrating (8%). khakiBee was alarmed to find that resubmitting the same prompt could generate different programs, commenting “You feel like you’ve made progress, and then because it did a different thing the next time, it’s like, what do I change? I’m trying to change what I give to the cow. And then that should change what the cow is doing. But if I’m not changing anything, why is that changing?” Some students also experienced the opposite issue: despite changing their descriptions, the model generated the same incorrect function repeatedly. purpleCarp commented, “Sometimes I changed my […] description and it just repeated the code the same. And it’s just very frustrating”. This highlights the difficulty of working with stochastic models: students expect the model output to be faithful to their descriptions.
@@ 7.2. Which Problems Do Students Say Are Hard?
We find that pass@1 and eventual success rates both vary by category (Table 4). We fit a binomial mixed-effects model to prompt success (1 if the prompt succeeded; 0 otherwise), with fixed effects of category, institution, and their interaction, and random effects of problem and participant (see Appendix, Table 12). A statistically reliable difference in success was observed only for Sorting problems, which were the most challenging ($p$=0.045). Participants from Oberlin struggled more in the Nested category compared to other students, but the effect is not statistically reliable ($p$=0.063).
@@ 7.4. What Do Students Do When They Encounter Unfamiliar Python?
Completion
@@
Question Is this code you would write yourself?
Student Responses Wellesley: Yes, Oberlin: No
Figure 6. An example code completion for the problem exp – this was generated by multiple different prompts. The completion was rated differently by Oberlin and Wellesley students, likely due to the list comprehension.